OpenStudy (anonymous):

Please help me! Will give medal :) Match each shape with its surface area. Use three point one four for pi and round to the nearest whole number where necessary.

OpenStudy (anonymous):

Match each term with definition. 1.square pyramid: base length 6 meters, slant height 10 meters 2.cone: radius 4 meters, slant height 8 meters 3.square pyramid: base length 5 meters, slant height 11 meters 4.cone: radius 6 meters, slant height 4 meters A)151m B)156m C)135m D)188m

OpenStudy (anonymous):

@ElisaNeedsHelp help me please

OpenStudy (elisaneedshelp):

I g2g for now bye sorry ill see when i get back

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

If I just tell you the formulas for calculating the surface areas, then can you just put the values and find the answers?

OpenStudy (anonymous):

umm maybe

mathslover (mathslover):

FOR CONE : Lateral Surface area = \(\pi \times r \times l\) ; l = slant height FOR Right Square Pyramid it is : \( = a^2 + 2a \sqrt{\cfrac{a^2}{4} + h}\) a= base edge h - height (NOT SLANT HEIGHT)

mathslover (mathslover):

If you are given with slant height in case of Right Square Pyramid then : SA = \(a^2 + 2\times a \times l \)

OpenStudy (anonymous):

ok thank you!

mathslover (mathslover):

You're welcome :) Good Luck and if you have any problem in working out with this , let me know.

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

@Shizen

OpenStudy (anonymous):

@shadow22 ... try using the formulas that @mathslover gave

OpenStudy (anonymous):

I tried for number 4. but the answer is wrong

OpenStudy (anonymous):

total surface area of cone = pi r ( r + l) where r = radius and l = slant height

OpenStudy (anonymous):

3.14*6*4=75.36 for #4

OpenStudy (anonymous):

no use the formula that i gave pi* radius ( radius + slant height)

OpenStudy (anonymous):

@shadow22 ?? there ?

OpenStudy (anonymous):

I got 117.04

OpenStudy (anonymous):

well lemme show you the work its the question ----> .cone: radius 6 meters, slant height 4 meters and formula for Total Surface Area of cone is pi *radius ( radius + slant height ) plugging in the values we get pi *6 ( 6 + 4) =>pi * 6 * 10 = now solve it further..!

OpenStudy (anonymous):

i got 188.4 now

OpenStudy (anonymous):

:D... and now round it up..

OpenStudy (anonymous):

188

OpenStudy (anonymous):

:D